### v3

@@ 6.3 Conclusive Summary
The studies reported in [79, 15, 16] are grounded in the empirical analysis of microservices systems to identify a multitude of issues, such as faults, bad smells, and performance issues faced by practitioners while designing, developing, and deploying microservices systems. To complement empiricism in microservices research and development, our proposed study mined the social coding platform (i.e., 15 open-source microservices systems on GitHub) and identified issues faced by developers with improvement and validation by microservices practitioners. While there is work on mining reusable knowledge [10, 67] - patterns and best practices – from software repositories to develop microservices systems, there is no research on mining knowledge, overseeing the broader microservices system development life cycle to streamline the plethora of issues, their causes and fixing strategies. Our study primarily focuses on survey-driven validation of the issues, causes, and solutions of microservices systems by practitioners, and complements the body of research comprising some of the recent industrial studies on evolvability [76], migration [40], and debugging [15] of microservices systems.
We provide a comparative analysis of the most relevant existing research and our proposed study in Table 7 based on (i) focus of investigation: specific aspects to be empirically investigated, (ii) research method: specific approaches employed for investigation, (iii) number of issues: specific number of the identified issues, and (iv) method for evaluation: specific method for evaluation. Additionally, the study reference, year of publication, and contribution of the research study are provided as complementary information for a fine-grained presentation of the results. For example, an empirical study reported in [15] published in 2021 relies on an industrial survey (i.e., applied research method) to investigate common faults and debugging practices in microservices systems (i.e., focus of investigation). The study results (i.e., main contribution) highlight typical faults, state-of-practice on debugging them, and indicate the challenges faced by developers for microservices systems.
Table 7: Comparative analysis and summary of existing research vs. this research
@@
[15] 2021 Fault analysis, Debugging Industry survey Not explicit mentioned
Contribution: An industrial survey to learn typical faults, current practice of debugging, and the challenges faced by developers for microservices systems.
[16] 2020 Performance issues MSA application monitoring Not explicit mentioned
Contribution: MicroRCA, as a solution to identify root causes of performance issues in MSAs by monitoring application performance and resource utilization.
[76] 2019 Evolvability analysis Structured interviews 16 types of challenges
Contribution: A qualitative study to explore the tools, metrics, patterns, and practitioners’ view on practices and challenges of microservices system evolvability.
[79] 2020 Code smells Static code analysis 11 types code smells
@@
[82] 2023 Security points Repository mining
@@
[83] 2022 Microservice problems and solutions Repository mining 14 problems, 47 solutions
@@
[84] 2022 Architectural smells Interviews, Repository mining 6 types of architectural smells
@@
[75] 2023 Antipattern specification and detection Repository mining, Static analysis Catalog of antipatterns
@@
[72] 2024 Dataset of OSS microservice projects Repository mining, Manual curation 378 projects
@@
[73] 2024 API evolution strategies, challenges, and automation Practitioner interviews (17), Proposed automation 6 strategies, 6 challenges
@@
[74] 2025 Dataset of microservice applications and workloads Repository mining, Empirical analysis 553 applications, 8 workloads
@@
Our previous study [18] 2022 Issues and Causes in microservices systems Repository mining 138 types of issues, 109 types of causes
@@
This study 2023 Issues, Causes, and Solutions in microservices systems Repository mining (2,641 issues), Structured interviews (15 interviews), Practitioners’ survey (150 responses) 402 types of issues, 228 types of causes, 177 types of solutions
@@ 7 Conclusions
This paper empirically investigates the issues, causes, and solutions of microservices systems by employing a mixed-methods approach. Our study collected data from 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub, 15 interviews and an online survey completed by 150 practitioners. The primary contribution of this work is rooted in the taxonomies of issues, causes, and solutions in microservices systems. The taxonomy of issues consists of 19 categories, 54 subcategories, and 402 types of issues. The taxonomy of causes contains 8 categories, 26 subcategories, and 228 types of causes, whereas the taxonomy of solutions includes 8 categories, 32 subcategories, and 177 types of solutions. Overviewing the results of this study, the major issues for microservices systems are Technical Debt, Continuous Integration and Delivery, and Exception Handling issues. The majority of the issues occur due to General Programming Errors, Missing Features and Artifacts, and Invalid Configuration and Communication. These issues are mainly addressed by Fixing, Adding, and Modifying artifacts.
Based on the results and our observations in this study, we proposed the future directions in three aspects:
New Techniques: Proposing (i) techniques for controlling TD through the design of microservices systems, (ii) multi-layered security solutions for fine-grained security management, (iii) architecting and design techniques for microservices systems with a particular focus on highly resilient and low coupled microservices systems, (ii) solutions to trace and isolate communication issues to increase fault tolerance, (iv) techniques for dynamically optimizing configuration settings of microservices, (v) techniques for identifying critical paths for improving performance and removing performance bottlenecks due to configuration issues, (vi) techniques for detecting security breach points during the configuration of microservices systems, (vii) strategies to automatically test APIs, load, and application security for microservices systems, (viii) techniques for dynamically assigning storage platforms according to the requirements of microservices, (ix) techniques for storing and accessing decentralized and shared data without losing the independence of individual microservices, (x) techniques for optimizing resource utilization (e.g., CPU and GPU usage), (xi) a framework for improving scalability to increase microservices performance, and (xii) techniques for converting code of one language to another used in microservices systems.
Empirically-Grounded Studies and Guidelines: Conducting empirical studies to (i) investigate technical debt at the code, design, and communication level of microservices systems, (ii) explore the efforts required to fix the build issues in microservices systems at different phases (e.g., compilation and linking phases), and (iii) compare the types of problems and the needed efforts to fix the build issues in microservices systems with monolithic systems. We also argue for preparing the strategies and guidelines for (i) addressing cd pipeline errors, (ii) establishing issue, cause, and solution knowledge base for microservices systems in the multi-cloud (e.g., AWS, Google Cloud) containerized environment, (iii) organizing polyglot databases for efficient read and write operations by considering performance, and (iv) addressing microservices security vulnerability and related risks at various levels, such as data centers, cloud providers, virtualization, communication, and orchestration.
Tools: Designing and developing intelligent tools (i) to identify, measure, prioritize, monitor, and prevent TD in microservices systems, (ii) for distributed tracing and real-time performance monitoring, and (iii) to fix Networking issues like localhost, ip address, and webhook errors. Moreover, The outcome of our study (e.g., issue, cause, and solution taxonomies) can help propose a systematic process and develop an intelligent recommendation system to (semi-)automatically identify the issues and causes of microservices system development, as well as recommend the solutions to address those issues. The recommendation system can assist microservices practitioners in efficiently and effectively designing, developing, and operationalizing microservices systems.
In conclusion, this paper has presented an overview of various issues that can occur in microservices systems. We have discussed the challenges and pitfalls in the design and implementation of microservices systems and highlighted the patterns and trends in the types of issues that arise. Additionally, we have provided insights into the solutions for addressing these issues. One other expectation that we can relate with our study is to not only contribute to the existing body of knowledge of issues, causes, and solutions in microservices systems — that lacks methodological and empirical rigor — but also to encourage other researchers to explore deeper into this highly important research area. Empirical knowledge of the nature of issues in microservices systems can help organizations to develop a better understanding of the challenges and opportunities that microservices architecture brings and how to address them.
Data availability
Link to our dataset is in the reference [19].

### v4

@@ 6.3 Conclusive Summary
The studies reported in [A2-michael2023empirical, A66-genfer2025understanding, A70-cerny2023catalog, A76-nogueira2024insights] are grounded in the empirical analysis of microservices systems to identify a multitude of issues, such as faults, bad smells, and performance issues faced by practitioners while designing, developing, and deploying microservices systems. To complement empiricism in microservices research and development, our proposed study mined the social coding platform (i.e., 15 open-source microservices systems on GitHub) and identified issues faced by developers with improvement and validation by microservices practitioners. While there is work on mining reusable knowledge [A10-daniel2023towards, A67-soldani2021mutosca] - patterns and best practices – from software repositories to develop microservices systems, there is no research on mining knowledge, overseeing the broader microservices system development life cycle to streamline the plethora of issues, their causes and fixing strategies. Our study primarily focuses on survey-driven validation of the issues, causes, and solutions of microservices systems by practitioners, and complements the body of research comprising some of the recent empirical studies on evolvability [A77-mohottige2025reengineering, A78-abdellatif2025identifying], migration [A40-cerny2025analyzing], and bad smells [A15-zhang2024trace] of microservices systems.
We provide a comparative analysis of the most relevant existing research and our proposed study in Table 9 based on (i) focus of investigation: specific aspects to be empirically investigated, (ii) research method: specific approaches employed for investigation, (iii) number of issues: specific number of the identified issues, and (iv) method for evaluation: specific method for evaluation. Additionally, the study reference, year of publication, and contribution of the research study are provided as complementary information for a fine-grained presentation of the results. For example, an empirical study reported in [A14-avritzer2025architecture] published in 2025 focuses on the detection of issues and anti-patterns of microservice performance by adopting metric-based analysis that investigates the call graphs between microservices. The main contribution of this study is to a tool-supported analysis of performance anti-patterns and issues in microservice systems.
Table 9: Comparative analysis and summary of existing vs. proposed research
@@
[A14-avritzer2025architecture] 2025 Detection of issues and anti-patterns of microservice performance Empirical study (metric analysis) Service call graph
Contribution: Tool-based assessment of Train Ticket benchmark, showing correlations between performance and architecture issues via service call graph analysis.
[A66-genfer2025understanding] 2025 Identification and evaluation of security tactics for microservice systems Controlled experiments with 70 participants to identify security features Security tactics applied in an MSA system
Contribution: Automatic generation of security models and the utilization of security-based metrics to guide software architectures through the assessment of security tactics employed within microservice systems.
[A10-daniel2023towards] 2023 Detection of patterns as best practices to architect microservice systems Empirical study (data mining) Architecture of open source system, i.e., MSA-based platform
Contribution: Detection of microservice patterns and two case studies with real-world applications to evaluate the accuracy and applicability of the proposed solution.
[C-2020Automated] 2020 Code smells Static code analysis 11 types code smells
@@
[nasab2022ess] 2023 Security points Repository mining
@@
[A15-zhang2024trace] 2024 Root-cause analysis and localisation of performance issues in microservice systems Controlled experiments based on sequential pattern mining and spectrum analysis of microservices Service interaction and coordination calls
Contribution: TraceContrast is a root cause localization tool that identifies bugs by comparing failed sequences against successful ones to pinpoint the specific combination of attributes causing the error.
[wu2022way] 2022 Microservice problems and solutions Repository mining 14 problems, 47 solutions
@@
[zhong2022impacts] 2022 Architectural smells Interviews, Repository mining 6 types of architectural smells
@@
[tighilt2023maintenance] 2023 Antipattern specification and detection Repository mining, Static analysis Catalog of antipatterns
@@
[amoroso2024dataset] 2024 Dataset of OSS microservice projects Repository mining, Manual curation 378 projects
@@
[lercher2024managing] 2024 API evolution strategies, challenges, and automation Practitioner interviews (17), Proposed automation 6 strategies, 6 challenges
@@
[lubas2025microservice] 2025 Dataset of microservice applications and workloads Repository mining, Empirical analysis 553 applications, 8 workloads
@@
Our previous study [waseem2021nature] 2022 Issues and Causes in microservices systems Repository mining 138 types of issues, 109 types of causes
@@
This study 2025 Issues, Causes, and Solutions in microservices systems Repository mining (2,641 issues), Structured interviews (15 interviews), Practitioners’ survey (150 responses) 402 types of issues, 228 types of causes, 177 types of solutions
@@ 7 Conclusions
This paper empirically investigates the issues, causes, and solutions of microservices systems by employing a mixed-methods approach, combining 2,641 GitHub issues from 15 open-source microservices systems, 15 interviews, and an online survey completed by 150 practitioners. The primary contribution of this work is the empirically grounded taxonomies of Issues, Causes, and Solutions in microservices systems, together with the mapping between them. Overviewing our results, the major Issue Categories are Technical Debt, Continuous Integration and Delivery, and Exception Handling, while the most common Cause Categories include General Programming Errors, Missing Features and Artifacts, and Invalid Configuration and Communication. These issues are mainly addressed through Fixing, Adding, and Modifying artifacts.
Finally, by linking Issue Categories to recurring Cause Categories and Solution Categories, our results provide an evidence base for both researchers and practitioners to understand where microservices problems concentrate and how they are typically addressed in OSS practice. Directions for future research are summarised in Section 4.3.
Data availability and ethics statement
The replication package (Ref. [replpack]) contains the study material and aggregated data used in this paper (e.g., survey questionnaire, interview and survey supporting artifacts). Participation in the interviews and survey was voluntary and consent-based. We applied data minimization: the instruments do not collect direct identifiers (e.g., names, email addresses, IP addresses, or organization names), and results are reported only in aggregated or anonymized form.